OpenStudy (anonymous):

Solve |3x-4|+5≤27 & how would I graph this??

OpenStudy (anonymous):

first get rid of the 5 on the left. \[\left| 3x-4 \right|\le 22\]Now deal with the absolute value. Recall the definition of absolute value. We have:\[(3x-4) \le 22\]or,\[-(3x-4)\le 22\]You have to solve each of these separately to get the full solution. Are you good from here?

OpenStudy (anonymous):

@eseidl Yep! I'm at 3x≤26 or 3x≤-18

OpenStudy (anonymous):

I get:\[x \le 8.67\](rounded), or\[x \ge -6\]remember to flip the inequality when you divide by the -3. Your solution on the number line is:Or,\[-6 \le x \le \frac{26}{3}\]